# Exploring the Frontier of Vision-Language Models: A Survey of Current Methodologies and Future Directions

### v1

@@ Abstract
The advent of Large Language Models (LLMs) has significantly reshaped the trajectory of the AI revolution. Nevertheless, these LLMs exhibit a notable limitation, as they are primarily adept at processing textual information. To address this constraint, researchers have endeavored to integrate visual capabilities with LLMs, resulting in the emergence of Vision-Language Models (VLMs). These advanced models are instrumental in tackling more intricate tasks such as image captioning and visual question answering. In our comprehensive survey paper, we delve into the key advancements within the realm of VLMs. Our classification organizes VLMs into three distinct categories: models dedicated to vision-language understanding, models that process multimodal inputs to generate unimodal (textual) outputs and models that both accept and produce multimodal inputs and outputs. This classification is based on their respective capabilities and functionalities in processing and generating various modalities of data.We meticulously dissect each model, offering an extensive analysis of its foundational architecture, training data sources, as well as its strengths and limitations wherever possible, providing readers with a comprehensive understanding of its essential components. We also analyzed the performance of VLMs in various benchmark datasets. By doing so, we aim to offer a nuanced understanding of the diverse landscape of VLMs. Additionally, we underscore potential avenues for future research in this dynamic domain, anticipating further breakthroughs and advancements.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has marked the onset of a transformative era in Artificial Intelligence, reshaping the entire landscape. Research labs, spanning both academia and industry, are actively engaged in a competitive race to advance the capabilities of LLMs. Yet, a notable limitation has come to the forefront – these models are confined to processing a singular modality of data, specifically text. This constraint underscores a pivotal challenge in the ongoing pursuit of refining LLMs to operate seamlessly across multiple modalities, marking a crucial avenue for further innovation in the realm of AI.
@@
In response to these limitations, researchers have pioneered a cutting-edge class of neural models known as Vision-Language Models (VLMs). These models intricately combine visual and textual information, showcasing a remarkable proficiency in comprehending and generating content that involves both images and text. Designed to excel in tasks such as image captioning, responding to visual queries, and generating images based on textual descriptions, VLMs exhibit a versatile capability. Their seamless integration of visual and linguistic modalities positions them at the forefront of technological advancements, allowing them to navigate the intricate interplay between images and text with unparalleled finesse.
In recent times, major research labs have been consistently introducing innovative VLMs, including DeepMind’s Flamingo, Salesforce’s BLIP, and OpenAI’s CLIP. Examples like GPT-4 (V) and Gemini showcase how chatbots are evolving within the realm of VLMs. Notably, not all multimodal models are VLMs; for instance, text-to-image models like Midjourney and DALL-E Ramesh et al. (2021) lack a language generation component, underscoring the diverse landscape of multimodal AI landscape. The general architecture of a VLM consists of an image and text encoder to generate the embeddings which are then fused in an image-text fusion layer and this fused vector is passed through an LLM to generate the final visually aware generated text. The working of a VLM is shown in Figure-2
@@
Recent surveys in this area like Wang et al. (2023b) mostly explore various pre-trained techniques and datasets used for developing multimodal models, Yin et al. (2023) explores various key techniques in training various multimodal language models. Wu et al. (2023a) provides practical applications and guidance in using multimodal language models. The most recent one by Zhang et al. (2024) covers around 26 most recent VLMs in depth. In contrast to previous surveys, none have systematically classified Vision-Language Models (VLMs) based on their input-processing and output-generation capabilities. Our survey addresses this gap by providing a thorough categorization of VLMs, shedding light on the intricacies of their functionalities. We extensively analyze the performance of different VLMs across benchmark datasets, notably including the latest MME benchmark, providing comprehensive insights. Our survey represents the most comprehensive and current compilation of VLMs to date, encompassing approximately 70 models. It serves as the ultimate guide for users navigating the evolving realm of vision-language models, offering the most current and comprehensive insights in this pioneering field of research.
@@ 2 Vision-Language Model (VLMs)
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=center, text centered, minimum width=5em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=15em,font=,, where level=2text width=14em,font=,, where level=3minimum width=10em,font=,, where level=4text width=26em,font=,, where level=5text width=20em,font=,, [ Vision-Language Models, for tree=fill=paired-light-red!70, text width=22em [ Vision-Language
@@
§2.1, for tree=fill=paired-light-yellow!45, text width=13em [ Handles only Images, for tree=fill=bg22, text width=13.5em [ CLIP Radford et al. (2021), for tree=fill=pg58,text width=30em ] [ AlphaCLIP sun2023alphaclip, for tree=fill=pg58,text width=30em ] [ MetaCLIP Xu et al. (2023b), for tree=fill=pg58,text width=30em ] [ GLIP Li et al. (2022a), for tree=fill=pg58, text width=30em ] [ ImageBind Girdhar et al. (2023), for tree=fill=pg58, text width=30em ] [ VLMo Bao et al. (2022), for tree=fill=pg58,text width=30em ] ] [ Handles Videos, for tree=fill=bg22, text width=13.5em [ VideoClip Xu et al. (2021), for tree=fill=pg58, text width=30em ] [ VideoMAE Tong et al. (2022), for tree=fill=pg58, text width=30em ] ] ] [ Text Generation with
@@
§2.2, for tree=fill=paired-light-yellow!45, text width=13em [ Handles only Images, for tree=fill=bg22, text width=13.5em [ GPT-4V , for tree=fill=bg16,text width=30em ] [ LLaVaLiu et al. (2023a) LLaVa-1.5Liu et al. (2023d)
LLaVa-PlusLiu et al. (2023e)
BakLLaVaLiu et al. (2023c) ,text width=30em, for tree=fill=bg16 ] [ Flamingo Alayrac et al. (2022),text width=30em, for tree=fill=bg16 ] [ IDEFICS Laurençon et al. (2023),text width=30em, for tree=fill=bg16 ] [ PaLI Chen et al. (2023f), text width=30em, for tree=fill=bg16 ] [ Qwen-VL Bai et al. (2023b), text width=30em, for tree=fill=bg16 ] [ Fuyu-8B Bavishi et al. (2023), text width=30em, for tree=fill=bg16 ] [ SPHINX Lin et al. (2023a), text width=30em, for tree=fill=bg16 ] [ Mirasol3B Piergiovanni et al. (2023), text width=30em, for tree=fill=bg16 ] [ MiniGPT-4 Zhu et al. (2023b) MiniGPT-v2 Chen et al. (2023d), text width=30em, for tree=fill=bg16 ] [ CogVLM Wang et al. (2023a), text width=30em,for tree=fill=bg16 ] [ Ferret You et al. (2023), text width=30em, for tree=fill=bg16 ] [ LaVIN Luo et al. (2023), text width=30em, for tree=fill=bg16 ] [ PALM-E Driess and et al (2023), text width=30em, for tree=fill=bg16 ] [ InstructBLIP Dai and et al. (2023)
BLIP Li et al. (2022b) BLIP-2 Li et al. (2023d), for tree=fill=bg16, text width=30em ] [ KOSMOS-1 Huang et al. (2023) KOSMOS-2 peng2023kosmos2, for tree=fill=bg16,text width=30em ] [ MultiInstruct Xu et al. (2022),for tree=fill=bg16,text width=30em ] [ Frozen Bain et al. (2021),for tree=fill=bg16,text width=30em ] [ CoVLM Li et al. (2023b),for tree=fill=bg16,text width=30em ] [ BEiT-3 Wang et al. (2022),for tree=fill=bg16, text width=30em ] [ mPLUG-2 Xu et al. (2023a),for tree=fill=bg16,text width=30em ] [ X2-VLM Zeng et al. (2023),for tree=fill=bg16,text width=30em ] [ Lyrics Lu et al. (2023),for tree=fill=bg16,text width=30em ] [ Prismer Liu et al. (2023b),for tree=fill=bg16,text width=30em ] [ X-FM Zhang et al. (2023b),for tree=fill=bg16,text width=30em ] [ MM-REACT Yang et al. (2023b),for tree=fill=bg16,text width=30em ] [ PICa Yang et al. (2022),for tree=fill=bg16,text width=30em ] [ PNP-VQA Tiong et al. (2022),for tree=fill=bg16,text width=30em ] [ Img2LLM Guo et al. (2023),for tree=fill=bg16,text width=30em ] [ SimVLM Wang et al. (2021),for tree=fill=bg16,text width=30em ] [ TinyGPT-V Yuan et al. (2023b),for tree=fill=bg16,text width=30em ] [ GPT4Tools Yang et al. (2023a),for tree=fill=bg16,text width=30em ] [ mPLUG-Owl Ye et al. (2023b), for tree=fill=bg16,text width=30em ] [ Ying-VLM Li et al. (2023e), for tree=fill=bg16,text width=30em ] ] [ Handles Videos, for tree=fill=bg22, text width=13.5em [LLaMA-VID Li et al. (2023c), text width=30em, for tree=fill=bg8] [ Video-LLaMA Zhang et al. (2023c), text width=30em, for tree=fill=bg8 ] [ ChatBridge Zhao et al. (2023b), text width=30em, for tree=fill=bg8 ] [ Video-ChatGPT Maaz et al. (2023), text width=30em, for tree=fill=bg8 ] [ VideoCoCa Yan et al. (2023), text width=30em, for tree=fill=bg8 ] [ VALOR Chen et al. (2023a),for tree=fill=bg8,text width=30em ] [ MACAW-LLM Lyu et al. (2023),for tree=fill=bg8,text width=30em ] [ PandaGPT Su et al. (2023),for tree=fill=bg8, text width=30em ] ] ] [ Multimodal Output with
@@
§2.3, for tree=fill=paired-light-yellow!45, text width=28em [ CoDiTang et al. (2023a) CoDi-2Tang et al. (2023b), text width=30em, for tree=fill=bg33 ] [ Gemini, text width=30em, for tree=fill=bg33 ] [ NeXT-GPT Wu et al. (2023c), text width=30em, for tree=fill=bg33 ] [ Video-Poet Kondratyuk et al. (2023), text width=30em, for tree=fill=bg33 ] ] ]
@@
Component Trainable Frozen LLM
@@
MiniGPT-V2 Chen et al. (2023d) Fuyu Bavishi et al. (2023), Qwen-VL Bai et al. (2023b)
@@
Qwen-VL Bai et al. (2023b) MiniGPT-V2 Chen et al. (2023d)
Image-Text Fusion: Some prominent examples are: Qformer: BLIP-2 Li et al. (2023d) Perceiver Resampler: Flamingo Alayrac et al. (2022) Full Connected Layers (MLP): Llava Liu et al. (2023c)
@@
In this section, we conduct a comprehensive examination of VLMs, categorizing them into three overarching classifications:
@@
  • Multimodal Output with Multimodal Input: This category delves into models that exhibit proficiency in generating multimodal outputs by processing multimodal inputs. This involves the synthesis of diverse modalities, such as visual and textual elements, to produce comprehensive and coherent results.We have shown this broad classification in the figure-1.222We have shown the most well-known VLMs in this taxonomy due to space constraints.
@@ 2.1 Vision-Language Understanding
CLIP Radford et al. (2021): CLIP, introduced by OpenAI, is a neural network proficient in grasping visual concepts through natural language guidance. It seamlessly identifies visual categories on diverse benchmarks, mirroring the "zero-shot" capabilities seen in GPT-powered models.By scaling a basic contrastive pre-training task, it achieved competitive zero-shot performance on diverse image classification datasets. CLIP gives a much more robust performance than the fine-tuned deep learning vision models in general for classification tasks. Though CLIP excels at common object recognition but struggles with abstract tasks, fine-grained classification, generalization, and wording sensitivity.
@@
AlphaCLIP Sun et al. (2023b): The model is an upgraded version of CLIP, incorporating an alpha channel for attentive region indication, enhancing awareness.Alpha-CLIP focuses on specific areas while maintaining CLIP’s recognition accuracy using an extended pipeline. It serves as a vision backbone for various applications, excelling in focused region attention but facing challenges with multiple objects and attention amplitude specification.
MetaCLIP Xu et al. (2023b): CLIP’s success lies in its data not just the model or pre-training objective. To address limited data transparency MetaCLIP was introduced which refines a raw data pool using metadata from CLIP’s concepts, outperforming CLIP on benchmarks with 400M image-text pairs from CommonCrawl.
@@
GLIP Li et al. (2022a): Inspired by CLIP, GLIP employs contrastive pre-training for language-image representations, emphasizing object-level alignment through phrase grounding. It redefines object detection as a vision-language task, utilizing deep fusion for improved representations. GLIP’s scalable pre-training on semantic-rich data enables automatic grounding box generation and robust zero/few-shot transferability to outperform baselines like CLIP on image captioning tasks and compete with a fully supervised dynamic head in downstream object detection tasks.
@@
VLMO Bao et al. (2022): Vlmo concurrently learns a dual encoder and a fusion encoder utilizing a modular Transformer network. The model adopts a Mixture-of-Modality-Experts (MOME) Transformer, comprising modality-specific experts and a shared self-attention layer in each block, providing notable modeling flexibility. Vlmo can be adeptly fine-tuned as a fusion encoder for vision-language classification tasks or employed as a dual encoder for efficient image-text retrieval, highlighting its versatility. The proposed stage-wise pre-training strategy maximizes the utilization of large-scale image-only and text-only data alongside image-text pairs, yielding state-of-the-art performance on diverse vision-language tasks such as visual question answering(VQA) and image retrieval.
@@
ImageBind Girdhar et al. (2023): The model learns a shared representation space by aligning various modalities’ embeddings to image embeddings, utilizing diverse paired data sources. It enables zero-shot recognition across modalities, leveraging web-scale image-text data and large-scale vision-language models like CLIP. This approach requires minimal training for application across different tasks and modalities. ImageBind utilizes large-scale image-text pairs and naturally paired self-supervised data across diverse modalities (audio, depth, thermal, IMU) to achieve strong emergent zero-shot classification and retrieval performance. It outperforms specialist models on audio benchmarks and demonstrates versatility in compositional tasks. Further enhancements include richer alignment data and adapting embeddings for specific tasks.
@@
VideoCLIP Xu et al. (2021): VideoCLIP focuses on pre-training a unified model for zero-shot video and text understanding without relying on labels for downstream tasks. It employs a contrastive learning framework with hard-retrieved negatives and overlapping positives for video-text pre-training. Key innovations include loosely temporally overlapping positive pairs and a retrieval-based sampling technique for negative pairs. The method utilizes contrastive loss and incorporates overlapped video-text clips for improved association.It is evaluated on various end tasks, showcasing state-of-the-art performance on video language datasets like Youcook2Zhou et al. (2018). The approach demonstrates significant advancements in zero-shot video-text understanding, outperforming previous work and even supervised approaches in some cases.
@@
VideoMAE Tong et al. (2022): VideoMAE is a self-supervised video pre-training method challenging the need for large-scale datasets. It adapts the masked autoencoder framework with a unique video tube masking strategy, achieving data efficiency on small datasets (3k-4k videos). It employs a Vision Transformer with joint space-time attention, demonstrating superior efficiency and effectiveness compared to traditional methods. VideoMAE outperforms in downstream tasks like action detection and holds potential for improvements through dataset expansion and integration of additional data streams. The paper acknowledges energy consumption concerns during pre-training but emphasizes VideoMAE’s practical value in scenarios with limited data availability.

### v2

@@ Abstract
The emergence of Large Language Models (LLMs) has profoundly altered the course of the AI revolution. Nevertheless, these LLMs exhibit a notable limitation, as they are primarily adept at processing textual information. To address this constraint, researchers have endeavored to integrate visual capabilities with LLMs, resulting in the emergence of Vision-Language Models (VLMs).These sophisticated models play a crucial role in addressing complex tasks like generating captions for images and responding to visual questions. In our comprehensive survey paper, we delve into the key advancements within the realm of VLMs. Our classification organizes VLMs into three distinct categories: models dedicated to vision-language understanding, models that process multimodal inputs to generate unimodal (textual) outputs and models that both accept and produce multimodal inputs and outputs. This classification is based on their respective capabilities and functionalities in processing and generating various modalities of data.We meticulously dissect each model, offering an extensive analysis of its foundational architecture, training data sources, as well as its strengths and limitations wherever possible, providing readers with a comprehensive understanding of its essential components. We also analyzed the performance of VLMs in various benchmark datasets. By doing so, we aim to offer a nuanced understanding of the diverse landscape of VLMs. Additionally, we underscore potential avenues for future research in this dynamic domain, anticipating further breakthroughs and advancements.
keywords:
Visual Language Models , Multimodal Language Models , Large Language Models , Generative-AI , Benchmark Datasets
journal: Journal of Knowledge Based System
@@ 1 Introduction
The rise of Large Language Models (LLMs) signifies the dawn of a transformative period in Artificial Intelligence, fundamentally restructuring the entire domain. Research labs, spanning both academia and industry, are actively engaged in a competitive race to advance the capabilities of LLMs. Yet, a notable limitation has come to the forefront – these models are confined to processing a singular modality of data, specifically text. This constraint underscores a pivotal challenge in the ongoing pursuit of refining LLMs to operate seamlessly across multiple modalities, marking a crucial avenue for further innovation in the realm of AI.
@@
In response to these limitations, researchers have pioneered a cutting-edge class of neural models known as Vision-Language Models (VLMs). These models intricately combine visual and textual information, showcasing a remarkable proficiency in comprehending and generating content that involves both images and text. Designed to excel in tasks such as responding to visual queries, image captioning and generating images based on textual descriptions, VLMs exhibit a versatile capability. Their seamless integration of visual and linguistic modalities positions them at the forefront of technological advancements, allowing them to navigate the intricate interplay between images and text with unparalleled finesse.
In recent times, major research labs have been consistently introducing innovative VLMs, including DeepMind’s Flamingo, Salesforce’s BLIP, and OpenAI’s CLIP. Examples like GPT-4 (V) and Gemini showcase how chatbots are evolving within the realm of VLMs. Notably, not all multimodal models are VLMs; for instance, text-to-image models like Midjourney and DALL-E [71] lack a language generation component, underscoring the diverse landscape of multimodal AI landscape. The general architecture of a VLM consists of an image and text encoder to generate the embeddings which are then fused in an image-text fusion layer and this fused vector is passed through an LLM to generate the final visually aware generated text. The working of a VLM is shown in Figure-2
@@
Recent surveys in this area like [87] mostly explore various pre-trained techniques and datasets used for developing multimodal models, [103] explores various key techniques in training various multimodal language models. [89] provides practical applications and guidance in using multimodal language models. The most recent one by [111] covers around 26 most recent VLMs in depth. In contrast to previous surveys, none have systematically classified Vision-Language Models (VLMs) based on their input-processing and output-generation capabilities. Our survey addresses this gap by providing a thorough categorization of VLMs, shedding light on the intricacies of their functionalities. We extensively analyze the performance of different VLMs across benchmark datasets, notably including the latest MME benchmark, providing comprehensive insights. Our survey represents the most comprehensive and current compilation of VLMs to date, encompassing approximately 70 models. It serves as the ultimate guide for users navigating the evolving realm of vision-language models, offering the most current and comprehensive insights in this pioneering field of research.
@@ 2 Vision-Language Model (VLMs)
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=center, text centered, minimum width=5em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=15em,font=,, where level=2text width=14em,font=,, where level=3minimum width=10em,font=,, where level=4text width=26em,font=,, where level=5text width=20em,font=,, [ Vision-Language Models, for tree=fill=paired-light-red!70, text width=22em [ Vision-Language
@@
§2.1, for tree=fill=paired-light-yellow!45, text width=13em [ Handles only Images, for tree=fill=bg22, text width=13.5em [ CLIP [70], for tree=fill=pg58,text width=30em ] [ AlphaCLIP [77], for tree=fill=pg58,text width=30em ] [ MetaCLIP [94], for tree=fill=pg58,text width=30em ] [ GLIP [41], for tree=fill=pg58, text width=30em ] [ ImageBind [29], for tree=fill=pg58, text width=30em ] [ VLMo [8], for tree=fill=pg58,text width=30em ] ] [ Handles Videos, for tree=fill=bg22, text width=13.5em [ VideoClip [93], for tree=fill=pg58, text width=30em ] [ VideoMAE [81], for tree=fill=pg58, text width=30em ] ] ] [ Text Generation with
@@
§2.2, for tree=fill=paired-light-yellow!45, text width=13em [ Handles only Images, for tree=fill=bg22, text width=13.5em [ GPT-4V , for tree=fill=bg16,text width=30em ] [ LLaVa[51] LLaVa-1.5[54]
LLaVa-Plus[55]
BakLLaVa[53] ,text width=30em, for tree=fill=bg16 ] [ Flamingo [4],text width=30em, for tree=fill=bg16 ] [ IDEFICS [40],text width=30em, for tree=fill=bg16 ] [ PaLI [17], text width=30em, for tree=fill=bg16 ] [ Qwen-VL [6], text width=30em, for tree=fill=bg16 ] [ Fuyu-8B [9], text width=30em, for tree=fill=bg16 ] [ SPHINX [48], text width=30em, for tree=fill=bg16 ] [ Mirasol3B [68], text width=30em, for tree=fill=bg16 ] [ MiniGPT-4 [120] MiniGPT-v2 [15], text width=30em, for tree=fill=bg16 ] [ CogVLM [85], text width=30em,for tree=fill=bg16 ] [ Ferret [104], text width=30em, for tree=fill=bg16 ] [ LaVIN [58], text width=30em, for tree=fill=bg16 ] [ PALM-E [26], text width=30em, for tree=fill=bg16 ] [ InstructBLIP [22]
BLIP [45] BLIP-2 [46], for tree=fill=bg16, text width=30em ] [ KOSMOS-1 [35] KOSMOS-2 [66], for tree=fill=bg16,text width=30em ] [ MultiInstruct [95],for tree=fill=bg16,text width=30em ] [ Frozen [7],for tree=fill=bg16,text width=30em ] [ CoVLM [43],for tree=fill=bg16,text width=30em ] [ BEiT-3 [86],for tree=fill=bg16, text width=30em ] [ mPLUG-2 [92],for tree=fill=bg16,text width=30em ] [ X2-VLM [108],for tree=fill=bg16,text width=30em ] [ Lyrics [56],for tree=fill=bg16,text width=30em ] [ Prismer [52],for tree=fill=bg16,text width=30em ] [ X-FM [110],for tree=fill=bg16,text width=30em ] [ MM-REACT [99],for tree=fill=bg16,text width=30em ] [ PICa [100],for tree=fill=bg16,text width=30em ] [ PNP-VQA [80],for tree=fill=bg16,text width=30em ] [ Img2LLM [33],for tree=fill=bg16,text width=30em ] [ SimVLM [88],for tree=fill=bg16,text width=30em ] [ TinyGPT-V [107],for tree=fill=bg16,text width=30em ] [ GPT4Tools [98],for tree=fill=bg16,text width=30em ] [ mPLUG-Owl [102], for tree=fill=bg16,text width=30em ] [ Ying-VLM [47], for tree=fill=bg16,text width=30em ] ] [ Handles Videos, for tree=fill=bg22, text width=13.5em [LLaMA-VID [44], text width=30em, for tree=fill=bg8] [ Video-LLaMA [112], text width=30em, for tree=fill=bg8 ] [ ChatBridge [114], text width=30em, for tree=fill=bg8 ] [ Video-ChatGPT [60], text width=30em, for tree=fill=bg8 ] [ VideoCoCa [97], text width=30em, for tree=fill=bg8 ] [ VALOR [11],for tree=fill=bg8,text width=30em ] [ MACAW-LLM [59],for tree=fill=bg8,text width=30em ] [ PandaGPT [74],for tree=fill=bg8, text width=30em ] ] ] [ Multimodal Output with
@@
§2.3, for tree=fill=paired-light-yellow!45, text width=28em [ CoDi[78] CoDi-2[79], text width=30em, for tree=fill=bg33 ] [ Gemini, text width=30em, for tree=fill=bg33 ] [ NeXT-GPT [91], text width=30em, for tree=fill=bg33 ] [ Video-Poet [38], text width=30em, for tree=fill=bg33 ] ] ]
@@
Component Trainable Frozen
LLM
@@
MiniGPT-V2 [15] Fuyu [9], Qwen-VL [6]
@@
Qwen-VL [6] MiniGPT-V2 [15]
Image-Text Fusion:
Some prominent examples are:
  • Qformer: BLIP-2 [46]
  • Perceiver Resampler: Flamingo [4]
  • Full Connected Layers (MLP): Llava [53]
@@
This section delves deeply into VLMs, organizing them into three broad classifications for a comprehensive review:
@@
  • Multimodal Output with Multimodal Input: This category delves into models that exhibit proficiency in generating multimodal outputs by processing multimodal inputs. This involves the synthesis of diverse modalities, such as visual and textual elements, to produce comprehensive and coherent results.We have shown this broad classification in the figure-1.111We have shown the most well-known VLMs in this taxonomy due to space constraints.
@@ 2.1 Vision-Language Understanding
CLIP [70]: CLIP, introduced by OpenAI, is a neural network proficient in grasping visual concepts through natural language guidance. It seamlessly identifies visual categories on diverse benchmarks, mirroring the ”zero-shot” capabilities seen in GPT-powered models.By scaling a basic contrastive pre-training task, it achieved competitive zero-shot performance on diverse image classification datasets. CLIP gives a much more robust performance than the fine-tuned deep learning vision models in general for classification tasks. Though CLIP excels at common object recognition but struggles with abstract tasks, fine-grained classification, generalization, and wording sensitivity.
@@
AlphaCLIP [76]: The model is an upgraded version of CLIP, incorporating an alpha channel for attentive region indication, enhancing awareness.Alpha-CLIP focuses on specific areas while maintaining CLIP’s recognition accuracy using an extended pipeline. It serves as a vision backbone for various applications, excelling in focused region attention but facing challenges with multiple objects and attention amplitude specification.
MetaCLIP [94]: The success of CLIP is attributed not only to its model or pre-training objective but also to its rich dataset. Introducing MetaCLIP, an innovative approach, aimed to overcome data transparency issues by refining raw data through metadata sourced from CLIP’s concepts. This enhancement demonstrated superior performance over CLIP on benchmarks utilizing a vast collection of 400M image-text pairs sourced from CommonCrawl.
@@
GLIP [41]: Inspired by CLIP, GLIP employs contrastive pre-training for language-image representations, emphasizing object-level alignment through phrase grounding. It redefines object detection as a vision-language task, utilizing deep fusion for improved representations. GLIP’s scalable pre-training on semantic-rich data enables automatic grounding box generation and robust zero/few-shot transferability to outperform baselines like CLIP on image captioning tasks and compete with a fully supervised dynamic head in downstream object detection tasks.
@@
VLMo [8]:Employing a modular Transformer architecture, VLMo concurrently acquires proficiency in both a dual encoder and a fusion encoder. It adopts a Mixture-of-Modality-Experts (MOME) Transformer framework, which integrates modality-specific experts alongside a shared self-attention mechanism within each module. This design affords substantial modeling flexibility. Vlmo exhibits noteworthy adaptability, functioning adeptly as a fusion encoder for tasks involving vision-language classification and as a dual encoder for efficient image-text retrieval. Its stage-wise pre-training approach optimally utilizes large-scale datasets encompassing images, texts, and their pairs. Consequently, VLMo achieves state-of-the-art performance across diverse vision-language tasks, including visual question answering (VQA) and image retrieval.
@@
ImageBind [29]: Imagebind learns a shared representation space by aligning embeddings from various modalities with image embeddings through diverse paired data sources. This facilitates zero-shot recognition across modalities, leveraging extensive web-scale image-text data and large-scale vision-language models such as CLIP. With minimal training required for deployment across different tasks and modalities, this approach proves highly adaptable. ImageBind utilizes a wealth of large-scale image-text pairs and naturally paired self-supervised data spanning multiple modalities (audio, depth, thermal, IMU) to achieve robust emergent zero-shot classification and retrieval capabilities. It surpasses specialized models on audio benchmarks and demonstrates versatility in compositional tasks. Additional enhancements include the incorporation of richer alignment data and the adaptation of embeddings tailored to specific tasks.
@@
VideoCLIP [93]: VideoCLIP’s objective is to pre-train a unified model capable of comprehending both video and text in a zero-shot manner, without dependence on labels for downstream tasks. Its approach involves employing a contrastive learning framework, integrating hard-retrieved negatives and overlapping positives during the pre-training phase for video-text comprehension. Noteworthy innovations include the incorporation of loosely temporally overlapping positive pairs and the utilization of a retrieval-based sampling technique for negative pairs. By leveraging contrastive loss and integrating overlapped video-text clips, VideoCLIP aims to enhance the association between different modalities.It is evaluated on various end tasks, showcasing state-of-the-art performance on video language datasets like Youcook2[118]. The approach demonstrates significant advancements in zero-shot video-text understanding, outperforming previous work and even supervised approaches in some cases.
@@
VideoMAE [81]: VideoMAE is a self-supervised video pre-training method challenging the need for large-scale datasets. It adapts the masked autoencoder framework with a unique video tube masking strategy, achieving data efficiency on small datasets (3k-4k videos). It employs a Vision Transformer with joint space-time attention, demonstrating superior efficiency and effectiveness compared to traditional methods. VideoMAE outperforms in downstream tasks like action detection and holds potential for improvements through dataset expansion and integration of additional data streams. The paper acknowledges energy consumption concerns during pre-training but emphasizes VideoMAE’s practical value in scenarios with limited data availability.